# Constraining interacting dark energy models with black hole superradiance

### v1

@@ Abstract
The recent preference for a dynamical dark energy (DE) from the Dark Energy Spectroscopic Instrument seems to call for interactions between DE and dark matter (DM), either from direct DM-DE interaction or indirect interaction induced by modified gravity. Therefore, an independent probe for these kinds of DE-DM interactions would be appealing from observational aspects. In this paper, we propose the black hole superradiance as a novel astrophysical probe for field-theoretic interacting DE-DM models, providing complementary constraints independent of large-scale cosmological observations. The core principle is that the DE-DM interaction can alter the effective mass of the superradiant ultralight boson, thereby modifying its superradiant instability rate around spinning black holes. We explore this connection through two distinct scenarios: a model where the DE field mediates a dark fifth force within the DM sector, affecting the superradiance from DM particles; and a novel mechanism where the DE field itself becomes superradiant due to the effective mass enhancement induced by dense DM spikes around supermassive black holes. By applying a statistical framework to black hole observations in both scenarios, we derive constraints on the fundamental DE-DM coupling strength. Although the current constraints are rather loose due to small samples and inaccurate measurements, our work provides new astrophysical constraints on these interacting DE-DM scenarios and establishes a new synergy between black hole physics and cosmology for probing the fundamental nature of the dark sector.
1 Introduction
@@
Typically, investigations into IDE models focus on their cosmological implications. By confronting theoretical predictions with data from the Cosmic Microwave Background and large-scale structure, the primary goal has been to understand the role of dark sector interactions in the overall cosmic expansion history. Our work, however, emphasizes a complementary perspective: the interaction within the dark sector can also profoundly impact physics on much smaller, astrophysical scales. We suggest that black hole (BH) superradiance provides a unique observational window to probe such small-scale effects. BH superradiance is a phenomenon where ultralight bosonic particles, interacting with a spinning BH, can trigger a rotational instability [zeldo1, Damour:1976kh, Teukolsky:1974yv, Press:1972zz] (see Ref. [Brito:2015oca] for a comprehensive review). This instability leads to the exponential growth of the boson field into a macroscopic “cloud” by rapidly extracting angular momentum from the BH. The potential observational consequences of BH superradiance have been widely explored, particularly the gravitational wave emission originating from the bosonic clouds [Arvanitaki:2014wva, Brito:2017wnc, Ghosh:2018gaw, Guo:2022mpr, liuGravitationalLaserStimulated2024, Kang:2024trj], the signatures imprinted on binary systems [Baumann:2018vus, baumannGravitationalColliderPhysics2020, Baumann:2021fkf, Baumann:2022pkl, Berti:2019wnn, Zhang:2018kib, Zhang:2019eid, Ikeda:2020xvt, Boskovic:2024fga, Takahashi:2021yhy, Takahashi:2024fyq, Cao:2023fyv, Tong:2022bbl, Tomaselli:2023ysb, Tomaselli:2024dbw, Guo:2024iye, Peng:2025zca, Guo:2025pea], and the electromagnetic signals [rosaStimulatedAxionDecay2018, Ikeda:2018nhb, Spieksma:2023vwl, Chen:2021lvo, Chen:2022kzv, Bai:2025yxm].
More pertinent to this work, BH superradiance has emerged as a powerful tool for probing fundamental physics beyond the Standard Model, especially within the dark sector [Arvanitaki:2009fg, Arvanitaki:2010sy, britoGravitationalWaveSearches2017, Stott:2018opm, Fernandez:2019qbj, Davoudiasl:2019nlo, Ng:2020ruv, Saha:2022hcd, chengConstrainsAxionlikeParticle2023, Aswathi:2025nxa, LIGOScientific:2021rnv, Yuan:2022bem]. The core principle is that the observation of highly-spinning BHs places stringent constraints on the existence of ultralight bosons within a characteristic mass range $\sim(G_{N}M)^{-1}$, where $M$ is the BH mass, $G_{N}$ is the Newton’s constant (in natural units with $\hbar=c=1$). This makes superradiance an effective probe for new light particles near BHs. Such probing becomes particularly interesting in scenarios where the dark sector involves multiple interacting species [Fukuda:2019ewf, East:2022ppo, Zhu:2025enp, Lyu:2025lue] or is influenced by the cosmological background, such as the cosmic neutrino background [Lambiase:2025twn].
@@
The work is structured as follows: In Section 2, we briefly review the theoretical foundations of IDE models, distinguishing between phenomenological and field-theoretic approaches. In Section 3, we establish our primary constraint methodology, detailing the physics of BH superradiance and the statistical framework used to constrain model parameters. We then apply this methodology to our specific scenarios. In Section 4, we present the Model I, where the DE scalar field acts as a mediator for a dark fifth force within the DM sector, and derive the corresponding constraints. In Section 5, we construct and analyze the Model II, a more intricate scenario where the DE field itself exhibits a superradiant instability induced by a DM spike. Finally, we conclude this work and discuss possible future directions in Section 6. Throughout this work, we adopt natural units with $\hbar=c=1$.
2 Interacting dark energy model
@@
2.1 Parameterization approach
In many cosmological studies, the dark components are treated as two separate, perfect fluids that permeate the Universe. In the context of an IDE, the standard assumption of individual energy conservation for DM and DE is relaxed. Instead, a non-gravitational energy exchange between them is permitted. This interaction is phenomenologically characterized by an energy transfer rate, $Q$, which modifies the conservation equations for both dark sector components as [Wang:2016lxa]
$\dot{\rho}_{\text{dm}}+3H\rho_{\text{dm}}=Q,$ (2.1)
$\dot{\rho}_{\text{de}}+3H(1+w)\rho_{\text{de}}=-Q.$ (2.2)
@@ 2.1 Parameterization approach
Within the phenomenological framework, the interaction term $Q$ is parameterized based on general physical considerations rather than derived from a specific fundamental theory. Its functional form is generally constructed to be proportional to the energy densities of the dark components and the Hubble rate, $H$, for dimensional consistency. Common linear interaction kernels in the literature include $Q=3H\xi\rho_{\text{dm}}$, $Q=3H\xi\rho_{\text{de}}$, or a combination of both [DiValentino:2017iww, Lucca:2020zjb, Zhai:2023yny, Bernui:2023byc, vanderWesthuizen:2025vcb]. In these expressions, $\xi$ represents a dimensionless coupling constant that quantifies the strength of the interaction. This parameterized approach provides a convenient framework to study the cosmological consequences of various energy exchange scenarios and to constrain the strength of such interactions with observational data.
2.2 Field-theoretic model
@@
$S=\int\mathrm{d}^{4}x\sqrt{-g}\left[\frac{R}{16\pi G_{N}}+\mathcal{L}_{\phi}\left(\phi,\partial_{\mu}\phi\right)+\mathcal{L}_{\chi}\left(\chi,\partial_{\mu}\chi\right)+\mathcal{L}_{\mathrm{int}}(\phi,\chi)\right]+S_{\mathrm{M}}.$ (2.3)
@@
$\mathcal{L}_{\phi}\left(\phi,\nabla_{\mu}\phi\right)$ $=-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-V(\phi),$ (2.4)
$\mathcal{L}_{\chi}\left(\chi,\nabla_{\mu}\chi\right)$ $=-\frac{1}{2}\nabla_{\mu}\chi\nabla^{\mu}\chi-U(\chi),$ (2.5)
$\mathcal{L}_{\mathrm{int}}(\phi,\chi)$ $=-W(\phi,\chi)\,.$ (2.6)
@@
$\dot{\rho}_{\phi}+3H\left(\rho_{\phi}+P_{\phi}\right)$ $=Q_{\phi},$ (2.7)
$\dot{\rho}_{\chi}+3H\left(\rho_{\chi}+P_{\chi}\right)$ $=Q_{\chi},$ (2.8)
$\dot{\rho}_{\mathrm{M}}+3H\left(\rho_{\mathrm{M}}+P_{\mathrm{M}}\right)$ $=0\,.$ (2.9)
This field-theoretic procedure provides a direct link between the macroscopic cosmological evolution and the microscopic properties of the dark sectors, thereby establishing a physical origin for the phenomenological term $Q$ in Eq. (2.2). As detailed in Appendix A, if the individual energy densities are defined such that their sum equals the total energy density of the dark sector, the fluid continuity equations are recovered with interaction terms that satisfy $Q_{\phi}=-Q_{\chi}$. This successfully maps the microscopic field theory to the phenomenological fluid model. For the specific case where the interaction potential $W(\phi,\chi)$ is assigned entirely to the dark matter component, the energy transfer rate is explicitly found to be $Q_{\chi}=-Q_{\phi}=W_{,\phi}\dot{\phi}$.
The general two-scalar model presented above serves as an illustrative framework to demonstrate how a microscopic field theory can give rise to an effective interaction consistent with the phenomenological fluid description. Our work builds upon this field-theoretic approach by constructing and analyzing two specific models derived from fundamental Lagrangians, which will be detailed in Sections 4 and 5, respectively, before we delve into the constraints from BH superradiance in Section 3.
Our first model (Model I, Section 4) is a direct realization of the framework discussed here, employing a trilinear coupling ($W(\phi,\chi)\propto\phi\chi^{2}$) between minimally coupled scalar fields representing DE and DM. This represents a scenario of direct interaction mediated by a potential term.
Our second model (Model II, Section 5) explores a different mechanism where the interaction arises indirectly through a non-minimal coupling to gravity. In this scenario, the DE field $\Phi$ couples conformally to the DM sector via the action term $S_{\mathrm{DM}}(\Psi;\Omega^{2}(\Phi)g_{\mu\nu})$. This coupling modifies the dynamics of the DE field itself, making its effective properties dependent on the local DM environment, without a direct potential term. This non-minimal coupling provides an alternative physical basis for interaction within the dark sector. Both models illustrate how macroscopic cosmic evolution links to microscopic dark sector particle properties, laying the groundwork for the later superradiance analysis.
3 Constraints from black hole superradiance
@@
3.1 Superradiance and black hole evolution
The interaction between a rotating BH and an ultralight scalar field $\phi$ with mass $\mu$ can trigger instability of the wave equation $(\square-\mu^{2})\phi=0$. This superradiant instability arises due to the presence of growing modes for $\phi$, which can be identified as the imaginary part of the eigenvalues in the associated Schrödinger equation. The superradiance rate depends on $\mu$ and is approximately maximized when the dimensionless “gravitational coupling” of the atom-like bosonic states around a spinning BH [Baumann:2019eav],
$\alpha\equiv GM\mu\approx 0.08\left(\frac{M}{10\mathrm{M}_{\odot}}\right)\left(\frac{\mu}{10^{-12}\mathrm{eV}}\right),$ (3.1)
@@ 3.1 Superradiance and black hole evolution
$\frac{\alpha}{m}<\frac{\tilde{a}}{2\left(1+\sqrt{1-\tilde{a}^{2}}\right)}$ (3.2)
is fulfilled. Here, $\tilde{a}=J/M^{2}$ is the dimensionless spin parameter of BH. In the non-relativistic approximation, the real component of the angular frequency is [Baumann:2019eav]
$\omega^{R}_{nlm}=\mu\left(1-\frac{\alpha^{2}}{2n^{2}}-\frac{\alpha^{4}}{8n^{4}}+\frac{(2\ell-3n+1)\alpha^{4}}{n^{4}(\ell+1/2)}+\frac{2\tilde{a}m\alpha^{5}}{n^{3}\ell(\ell+1/2)(\ell+1)}+O\left(\alpha^{6}\right)\right)\,$, (3.3)
@@
$\omega^{I}_{nlm}$ $=2m_{\phi}r_{+}\left(m\Omega_{H}-\mu\right)\alpha^{4l+4}\mathcal{A}_{nl}\mathcal{X}_{nl},$ (3.4)
$\mathcal{A}_{nl}$ $=\frac{2^{4l+1}(2l+n+1)!}{(l+n+1)^{2l+4}n!}\left(\frac{l!}{(2l)!(2l+1)!}\right)^{2},$ (3.5)
$\mathcal{X}_{lm}$ $=\prod_{j=1}^{l}\left[j^{2}\left(1-a_{*}^{2}\right)+4r_{+}^{2}\left(m\Omega_{H}-m_{\phi}\right)^{2}\right]\,.$ (3.6)
@@
$N_{\mathrm{max}}\approx 10^{76}\left(\frac{1}{m}\right)\left(\frac{\Delta\tilde{a}}{0.1}\right)\left(\frac{M}{10\mathrm{M}_{\odot}}\right)^{2}\,.$ (3.9)
3.2 Constraint methodology
@@
A simplified criterion for the second condition can be formulated by comparing the superradiance timescale, $\tau_{\mathrm{SR}}=\Gamma_{n\ell m}^{-1}$, with the BH timescale. Since the occupation number of the cloud grows exponentially, the time required to reach the maximum number of particles, $N_{\mathrm{max}}$, is proportional to $\tau_{\mathrm{SR}}\ln N_{\mathrm{max}}$. For the spin-down process to be effective, this must be shorter than the BH lifetime, leading to the condition  [^1] It should be noted that the value of $\Delta\tilde{a}$ used to calculate $N_{\mathrm{max}}$ is typically chosen to be on the order of the observational uncertainty of the BH spin. Because the condition on $\tau_{\mathrm{SR}}$ depends only logarithmically on $N_{\mathrm{max}}$ (and thus on $\Delta\tilde{a}$), even an order-of-magnitude difference in $N_{\mathrm{max}}$ does not significantly alter the derived limits, making this timescale comparison a robust estimation method [Hoof:2024quk].
$\tau_{\mathrm{SR}}<\frac{\tau_{\mathrm{BH}}}{\ln N_{\mathrm{max}}}\,.$ (3.10)
@@
This constraint methodology, based on BH spin observations, is general and can be applied to various models featuring ultralight bosons. For instance, in typical axion models (potentially including self-interactions), the constraints would be placed on the axion mass $m_{a}$ and decay constant $f_{a}$ [britoGravitationalWaveSearches2017, Stott:2018opm, Fernandez:2019qbj, Davoudiasl:2019nlo, Ng:2020ruv, Saha:2022hcd, chengConstrainsAxionlikeParticle2023, Aswathi:2025nxa]. However, our focus in this work is to constrain the IDE models, where the relevant parameters governing the superradiance of the ultralight boson (be it DM or DE) differ due to the DE-DM coupling.
@@
Specifically, for a given IDE model characterized by a set of fundamental parameters $\boldsymbol{\alpha}$, the superradiance predicts a maximally possible spin, $\tilde{a}^{\mathrm{crit}}$, that a BH of mass $M$ can sustain over its lifetime $\tau_{\mathrm{BH}}$. This defines an “exclusion boundary” or Regge trajectory in the BH parameter space, denoted as $\tilde{a}^{\mathrm{crit}}(M;\tau_{\mathrm{BH}},\boldsymbol{\alpha})$ [Hoof:2024quk]. For our Model I, the relevant parameter space is $\boldsymbol{\alpha}=(m_{\chi},\beta)$, while for Model II, it simplifies into $\boldsymbol{\alpha}=(\beta)$. It is essential to note that when applying this method to our IDE models, we are currently neglecting the potential self-interactions of the superradiating field itself (analogous to the effects controlled by $f_{a}$ in axion models).
@@
The physical implication remains straightforward: if an astrophysical BH is observed with a spin $\tilde{a}_{\mathrm{obs}}$ that is confidently above this theoretical maximum ($\tilde{a}_{\mathrm{obs}}>\tilde{a}^{\mathrm{crit}}$), then the specific set of model parameters $\boldsymbol{\alpha}$ that predicted that boundary is ruled out [Hoof:2024quk]. Conversely, a BH observed with a spin below the boundary ($\tilde{a}_{\mathrm{obs}}\leq\tilde{a}^{\mathrm{crit}}$) is consistent with the model, as it might simply have formed with a lower initial spin [Hoof:2024quk].
3.3 Statistical framework

### v2

@@ Abstract
The recent preference for a dynamical dark energy (DE) from the Dark Energy Spectroscopic Instrument seems to call for interactions between DE and dark matter (DM), either from direct DE-DM interaction or indirect interaction induced by modified gravity. Therefore, an independent probe for these kinds of DE-DM interactions would be appealing from observational aspects. In this paper, we propose the black hole superradiance as a novel astrophysical probe for field-theoretic interacting DE-DM models, providing complementary constraints independent of large-scale cosmological observations. The core principle is that the DE-DM interaction can alter the effective mass of the superradiant ultralight boson, thereby modifying its superradiant instability rate around spinning black holes. We explore this connection through two distinct scenarios: a model where the DE field mediates a dark fifth force within the DM sector, affecting the superradiance from DM particles; and a novel mechanism where the DE field itself becomes superradiant due to the effective mass enhancement induced by dense DM spikes around supermassive black holes. By applying a statistical framework to black hole observations in both scenarios, we derive constraints on the fundamental DE-DM coupling strength. Although the current constraints are rather loose due to small samples and inaccurate measurements, our work provides new astrophysical constraints on these interacting DE-DM scenarios and establishes a new synergy between black hole physics and cosmology for probing the fundamental nature of the dark sector.
I Introduction
@@
Typically, investigations into IDE models focus on their cosmological implications. By confronting theoretical predictions with data from the Cosmic Microwave Background and large-scale structure, the primary goal has been to understand the role of dark sector interactions in the overall cosmic expansion history. Our work, however, emphasizes a complementary perspective: the interaction within the dark sector can also profoundly impact physics on much smaller, astrophysical scales. We suggest that black hole (BH) superradiance provides a unique observational window to probe such small-scale effects. BH superradiance is a phenomenon where ultralight bosonic particles, interacting with a spinning BH, can trigger a rotational instability [128, 45, 114, 100] (see Ref. [28] for a comprehensive review). This instability leads to the exponential growth of the boson field into a macroscopic “cloud” by rapidly extracting angular momentum from the BH. The potential observational consequences of BH superradiance have been widely explored, particularly the gravitational wave emission originating from the bosonic clouds [10, 30, 59, 1, 127, 65, 86, 73], the signatures imprinted on binary systems [20, 19, 17, 18, 23, 130, 131, 69, 25, 110, 111, 36, 117, 115, 116, 63, 64, 95, 66, 82, 75], and the electromagnetic signals [102, 70, 107, 41, 42, 14].
More pertinent to this work, BH superradiance has emerged as a powerful tool for probing fundamental physics beyond the Standard Model, especially within the dark sector [11, 12, 29, 109, 54, 46, 92, 105, 43, 126, 67, 37, 13]. The core principle is that the observation of highly-spinning BHs places stringent constraints on the existence of ultralight bosons within a characteristic mass range $\sim(G_{N}M)^{-1}$, where $M$ is the BH mass, $G_{N}$ is the Newton’s constant (in natural units with $\hbar=c=1$). This makes superradiance an effective probe for new light particles near BHs. Such probing becomes particularly interesting in scenarios where the dark sector involves multiple interacting species [57, 52, 132, 89] or is influenced by the cosmological background, such as the cosmic neutrino background [76].
@@
The work is structured as follows: In Section II, we briefly review the theoretical foundations of IDE models, distinguishing between phenomenological and field-theoretic approaches. In Section III, we establish our primary constraint methodology, detailing the physics of BH superradiance and the statistical framework used to constrain model parameters. We then apply this methodology to our specific scenarios. In Section IV, we present the Model I, where the DE scalar field acts as a mediator for a dark fifth force within the DM sector, and derive the corresponding constraints. In Section V, we construct and analyze the Model II, a more intricate scenario where the DE field itself exhibits a superradiant instability induced by a DM spike. Finally, we conclude this work and discuss possible future directions in Section VI. Throughout this work, we adopt natural units with $\hbar=c=1$.
II Interacting dark energy model
@@
II.1 Parameterization approach
In many cosmological studies, the dark components are treated as two separate, perfect fluids that permeate the Universe. In the context of an IDE, the standard assumption of individual energy conservation for DM and DE is relaxed. Instead, a non-gravitational energy exchange between them is permitted. This interaction is phenomenologically characterized by an energy transfer rate, $Q$, which modifies the conservation equations for both dark sector components as [121]
$\dot{\rho}_{\text{dm}}+3H\rho_{\text{dm}}=Q,$ (1)
$\dot{\rho}_{\text{de}}+3H(1+w)\rho_{\text{de}}=-Q.$ (2)
@@ II.1 Parameterization approach
Within the phenomenological framework, the interaction term $Q$ is parameterized based on general physical considerations rather than derived from a specific fundamental theory. Its functional form is generally constructed to be proportional to the energy densities of the dark components and the Hubble rate, $H$, for dimensional consistency. Common linear interaction kernels in the literature include $Q=3H\xi\rho_{\text{dm}}$, $Q=3H\xi\rho_{\text{de}}$, or a combination of both [48, 88, 129, 22, 119]. In these expressions, $\xi$ represents a dimensionless coupling constant that quantifies the strength of the interaction. This parameterized approach provides a convenient framework to study the cosmological consequences of various energy exchange scenarios and to constrain the strength of such interactions with observational data.
II.2 Field-theoretic model
@@
$S$ $=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\frac{R}{16\pi G_{N}}+\mathcal{L}_{\phi}\bigl(\phi,\partial_{\mu}\phi\bigr)\right$.
$\left.+\mathcal{L}_{\chi}\bigl(\chi,\partial_{\mu}\chi\bigr)+\mathcal{L}_{\mathrm{int}}(\phi,\chi)\right]+S_{\mathrm{M}}\,.$ (3)
@@
$\mathcal{L}_{\phi}\left(\phi,\nabla_{\mu}\phi\right)$ $=-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-V(\phi),$ (4)
$\mathcal{L}_{\chi}\left(\chi,\nabla_{\mu}\chi\right)$ $=-\frac{1}{2}\nabla_{\mu}\chi\nabla^{\mu}\chi-U(\chi),$ (5)
$\mathcal{L}_{\mathrm{int}}(\phi,\chi)$ $=-W(\phi,\chi)\,.$ (6)
@@
$\dot{\rho}_{\phi}+3H\left(\rho_{\phi}+P_{\phi}\right)$ $=Q_{\phi},$ (7)
$\dot{\rho}_{\chi}+3H\left(\rho_{\chi}+P_{\chi}\right)$ $=Q_{\chi},$ (8)
$\dot{\rho}_{\mathrm{M}}+3H\left(\rho_{\mathrm{M}}+P_{\mathrm{M}}\right)$ $=0\,.$ (9)
This field-theoretic procedure provides a direct link between the macroscopic cosmological evolution and the microscopic properties of the dark sectors, thereby establishing a physical origin for the phenomenological term $Q$ in Eq. (2) (see also [3, 2, 72, 71, 15]). As detailed in Appendix A, if the individual energy densities are defined such that their sum equals the total energy density of the dark sector, the fluid continuity equations are recovered with interaction terms that satisfy $Q_{\phi}=-Q_{\chi}$. For the specific case where the interaction potential $W(\phi,\chi)$ is assigned entirely to the dark matter component, the energy transfer rate is explicitly found to be $Q_{\chi}=-Q_{\phi}=W_{,\phi}\dot{\phi}$.
The general two-scalar model presented above serves as an illustrative framework to demonstrate how a microscopic field theory can give rise to an effective interaction consistent with the phenomenological fluid description. Our work builds upon this field-theoretic approach by constructing and analyzing two specific models derived from fundamental Lagrangians, which will be detailed in Sections IV and V, respectively. Before turning to these models, we delve into the constraints from BH superradiance in Section III.
Our first model (Model I, Section IV) is a direct realization of the framework discussed here, employing a trilinear coupling ($W(\phi,\chi)\propto\phi\chi^{2}$) between minimally coupled scalar fields representing DE and DM. This represents a scenario of direct interaction mediated by a potential term.
Our second model (Model II, Section V) explores a different mechanism where the interaction arises indirectly through a non-minimal coupling to gravity. In this scenario, the DE field $\Phi$ couples conformally to the DM sector via the action term $S_{\mathrm{DM}}(\Psi;\Omega^{2}(\Phi)g_{\mu\nu})$. This coupling modifies the dynamics of the DE field itself, making its effective properties dependent on the local DM environment, without a direct potential term. This non-minimal coupling provides an alternative physical basis for interaction within the dark sector. Both models illustrate how macroscopic cosmic evolution links to microscopic dark sector particle properties, laying the groundwork for the later superradiance analysis.
III Constraints from black hole superradiance
@@
III.1 Superradiance and black hole evolution
The interaction between a rotating BH and an ultralight scalar field $\phi$ with mass $\mu$ can trigger instability of the wave equation $(\square-\mu^{2})\phi=0$. This superradiant instability arises due to the presence of growing modes for $\phi$, which can be identified as the imaginary part of the eigenvalues in the associated Schrödinger equation. The superradiance rate depends on $\mu$ and is approximately maximized when the dimensionless “gravitational coupling” of the atom-like bosonic states around a spinning BH [21],
$\alpha\equiv GM\mu\approx 0.08\left(\frac{M}{10\mathrm{M}_{\odot}}\right)\left(\frac{\mu}{10^{-12}\mathrm{eV}}\right),$ (10)
@@ III.1 Superradiance and black hole evolution
$\frac{\alpha}{m}<\frac{\tilde{a}}{2\left(1+\sqrt{1-\tilde{a}^{2}}\right)}$ (11)
is fulfilled. Here, $\tilde{a}=J/M^{2}$ is the dimensionless spin parameter of BH. In the non-relativistic approximation, the real component of the angular frequency is [21]
$\omega^{R}_{n\ell m}=\mu\Bigg($ $1-\frac{\alpha^{2}}{2n^{2}}-\frac{\alpha^{4}}{8n^{4}}+\frac{(2\ell-3n+1)\alpha^{4}}{n^{4}(\ell+1/2)}$
$\quad+\frac{2\,\tilde{a}\,m\,\alpha^{5}}{n^{3}\,\ell(\ell+1/2)(\ell+1)}+O\bigl(\alpha^{6}\bigr)\Bigg)\,$, (12)
@@
$\omega^{I}_{nlm}$ $=2m_{\phi}r_{+}\left(m\Omega_{H}-\mu\right)\alpha^{4l+4}\mathcal{A}_{nl}\mathcal{X}_{nl},$ (13)
$\mathcal{A}_{nl}$ $=\frac{2^{4l+1}(2l+n+1)!}{(l+n+1)^{2l+4}n!}\left(\frac{l!}{(2l)!(2l+1)!}\right)^{2},$ (14)
$\mathcal{X}_{lm}$ $=\prod_{j=1}^{l}\left[j^{2}\left(1-a_{*}^{2}\right)+4r_{+}^{2}\left(m\Omega_{H}-m_{\phi}\right)^{2}\right]\,.$ (15)
@@
$N_{\mathrm{max}}\approx 10^{76}\left(\frac{1}{m}\right)\left(\frac{\Delta\tilde{a}}{0.1}\right)\left(\frac{M}{10\mathrm{M}_{\odot}}\right)^{2}\,.$ (18)
III.2 Constraint methodology
@@
A simplified criterion for the second condition can be formulated by comparing the superradiance timescale, $\tau_{\mathrm{SR}}=\Gamma_{n\ell m}^{-1}$, with the BH timescale. Since the occupation number of the cloud grows exponentially, the time required to reach the maximum number of particles, $N_{\mathrm{max}}$, is proportional to $\tau_{\mathrm{SR}}\ln N_{\mathrm{max}}$. For the spin-down process to be effective, this must be shorter than the BH lifetime, leading to the condition  [^1] It should be noted that the value of $\Delta\tilde{a}$ used to calculate $N_{\mathrm{max}}$ is typically chosen to be on the order of the observational uncertainty of the BH spin. Because the condition on $\tau_{\mathrm{SR}}$ depends only logarithmically on $N_{\mathrm{max}}$ (and thus on $\Delta\tilde{a}$), even an order-of-magnitude difference in $N_{\mathrm{max}}$ does not significantly alter the derived limits, making this timescale comparison a robust estimation method [67].
$\tau_{\mathrm{SR}}<\frac{\tau_{\mathrm{BH}}}{\ln N_{\mathrm{max}}}\,.$ (19)
@@
This constraint methodology, based on BH spin observations, is general and can be applied to various models featuring ultralight bosons. For instance, in typical axion models (potentially including self-interactions), the constraints would be placed on the axion mass $m_{a}$ and decay constant $f_{a}$ [11, 12, 29, 109, 54, 46, 92, 105, 43, 126, 67, 37, 13]. However, our focus in this work is to constrain the IDE models, where the relevant parameters governing the superradiance of the ultralight boson (be it DM or DE) differ due to the DE-DM coupling.
@@
Specifically, for a given IDE model characterized by a set of fundamental parameters $\boldsymbol{\alpha}$, the superradiance predicts a maximally possible spin, $\tilde{a}^{\mathrm{crit}}$, that a BH of mass $M$ can sustain over its lifetime $\tau_{\mathrm{BH}}$. This defines an “exclusion boundary” or Regge trajectory in the BH parameter space, denoted as $\tilde{a}^{\mathrm{crit}}(M;\tau_{\mathrm{BH}},\boldsymbol{\alpha})$ [67]. For our Model I, the relevant parameter space is $\boldsymbol{\alpha}=(m_{\chi},\beta)$, while for Model II, it simplifies into $\boldsymbol{\alpha}=(\beta)$. It is essential to note that when applying this method to our IDE models, we are currently neglecting the potential self-interactions of the superradiating field itself (analogous to the effects controlled by $f_{a}$ in axion models).
@@
The physical implication remains straightforward: if an astrophysical BH is observed with a spin $\tilde{a}_{\mathrm{obs}}$ that is confidently above this theoretical maximum ($\tilde{a}_{\mathrm{obs}}>\tilde{a}^{\mathrm{crit}}$), then the specific set of model parameters $\boldsymbol{\alpha}$ that predicted that boundary is ruled out [67]. Conversely, a BH observed with a spin below the boundary ($\tilde{a}_{\mathrm{obs}}\leq\tilde{a}^{\mathrm{crit}}$) is consistent with the model, as it might simply have formed with a lower initial spin [67].
III.3 Statistical framework